OpenStudy (anonymous):

How do you factor 27x^3+108x^2+144x+64

OpenStudy (anonymous):

look for common factors in the coefficients first to make your life easier

OpenStudy (anonymous):

oops there are none, nvm

OpenStudy (anonymous):

i guess you are supposed to recognize this as a perfect cube

OpenStudy (anonymous):

I have answer choices if anyone would like to know them?

OpenStudy (anonymous):

\((a+b)^3=a^3+3a^2b+3ab^2+b^3\) put \(a=3x,b=4\)

OpenStudy (anonymous):

therefore the answer would be (3x+4)^3 ?

OpenStudy (anonymous):

in other words it is \((3x+4)^3\)

OpenStudy (anonymous):

yes exactly!

OpenStudy (anonymous):

Thanks so muuch , you guys are a life saver. and its free.thanks sooo much

OpenStudy (anonymous):

yw invoice is in the mail

OpenStudy (anonymous):

What does that mean?